OpenStudy (anonymous):

x + y = 6 2x - y = 3 The system of equations has no solution (has one solution) correct? is coincident

OpenStudy (anonymous):

Lets first solve the question . Rearrange the first Eq. makes x = 6 - y . Then substitute it in eq.2 which makes 2 ( 6-y ) - y = 3 , thus 12 - 2y - y = 3 . Therefore -3y = -9 . y = -3/-9 Thus y = 3 . X = 6 - y = 3 . The equations have only one solution

OpenStudy (anonymous):

im right, thanks

OpenStudy (anonymous):

Yes thats right

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Do me best response ! :D

OpenStudy (anonymous):

5x + 2y = 5 3x - y = 14 Which of the following equations is not a result of solving the system of equations by addition? 11x = 19 11y = -55 (11x = 33) correct?

OpenStudy (anonymous):

Y = 3X - 14 . 5X + 6X - 28 = 5 . Thus 11x = 33 . The answer will be 11y = -55. Because the question asks for IS NOT A RESULT .

OpenStudy (anonymous):

ohhh. thanks.

OpenStudy (anonymous):

x + 2y = 5 3x + 5y = 14 Solve the system of equations. (3, 1) (7, -1) (2, 3/2) is this correct?

OpenStudy (anonymous):

I had a little trouble with this one

OpenStudy (anonymous):

x = 5 - 2y . 3 ( 5 - 2y ) + 5y = 14 15 - 6y +5y = 14 15 - y = 14 15 - 14 = y y = 1

OpenStudy (anonymous):

x=y=3 : is only one solution if x,y belong to integer.

OpenStudy (anonymous):

x = 5 - 2 (1) x = 3

OpenStudy (anonymous):

Thanks. 7x + 2y = 4 y = x + 1 Solve the system of equations. (1/3, 4/3) correct? (2/9, 11/9) no solution

OpenStudy (anonymous):

7x + 2x + 2 = 4 9x = 2 x = 2/9 y = 2/9 + 1 = 11/ 9 so 1 solution

OpenStudy (anonymous):

i just need help with two more

OpenStudy (anonymous):

One number is 5 more than twice another and their sum is 23. Which of the following system of equations represents the word problem? x + y = 23 and y = 2x + 5 correct? y = x + 5 and y + 2(x + 5) = 23 y = 2(x + 5) and x + y = 23

OpenStudy (anonymous):

x + y = 23 . AND y = 2x + 5 x = 23 - y y = 2(23 - y ) + 5 y = 46 - 2y + 5 3 y = 51 y = 17 x = 23 - y x = 23 - 17 x = 6

OpenStudy (anonymous):

Any more question ?